OpenStudy (anonymous):

Ted's favorite coffee shop provides internet access to its customers for initial fee of $3 plus 20 cents per minute. A. How long can he use the internet if his budget is $15? B. Write an equation in slop-intercept form that represents the relationship between number of minutes,x and the cost of using wireless internet, y. C. Graph the linear function.

OpenStudy (anonymous):

3+20(m) is the formula u would use so 15-3=12 so u have 12$ left follow me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then you do $0.20(m) and try to get as close to 12.00 without going over follow me still?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

.20(59 minutes) is $11.80 so 59 minutes is the answer for a.

OpenStudy (anonymous):

B. y=.20x+3 do you see how i got this?

OpenStudy (anonymous):

yes i got the same thing

OpenStudy (anonymous):

okay so now for c. i may have to graph this out and send you the link is that okay?

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

.20x +3 is less than or equal to 15

OpenStudy (anonymous):

i was getting to that

OpenStudy (mertsj):

slope is 1/5

OpenStudy (anonymous):

y=1/5x+12 ok thats the formula for the graph

OpenStudy (mertsj):

\[.20x + 3\le 15\]

OpenStudy (anonymous):

what is that one for

OpenStudy (mertsj):

the cost is .20x + 3. It must be less than or equal to 15

OpenStudy (mertsj):

He can use the internet for one hour before he exceeds $15

OpenStudy (anonymous):

@mertsj you subtracted 3 form 15 so it would have to be \[\le 12\]

OpenStudy (mertsj):

What you want to do is graph y = 1/5x + 3. If y is 15, he has used up all his money. And at a y value of 15, x will be 60

OpenStudy (mertsj):